Please solve and show the work for this problem below. This is how the problem was listed exactly.

Lowe's Companies, Inc. Common Size Statement of Earnings 12 Months Ended in Millions $ Percentages
Feb. 01, 2019 Feb. 02, 2018 Feb. 03, 2017 Feb. 01, 2019 Feb. 02, 2018 Feb. 03, 2017
Net sales $ 71,309 $ 68,619 $ 65,017
Cost of sales 48,401 46,185 43,343
Gross margin 22,908 22,434 21,674
Expenses:
Selling, general and administrative 17,413 14,444 14,375
Depreciation and amortization 1,477 1,404 1,453
Operating income 4,018 6,586 5,846
Interest - net 624 633 645
Loss on extinguishment of debt 0 464 0
Pre-tax earnings 3,394 5,489 5,201
Income tax provision 1,080 2,042 2,108
Net earnings $ 2,314 $ 3,447 $ 3,093
Answer #1

For making the common size statement of earnings we will take the base for each year as its net sales amount. So for example the line item of selling, general and administrative expense for 2019 = 17,413/71,309 = 24.42%. The line item depreciation and amortization for 2018 will be = 1404/68619 = 2.05%. The line item interest for 2017 = 645/65017 = 0.99%

The results are shown below:

2019 2018 2017
Net sales 100.00% 100.00% 100.00%
Cost of sales 67.88% 67.31% 66.66%
Gross margin 32.12% 32.69% 33.34%
Expenses:
Selling, general and administrative 24.42% 21.05% 22.11%
Depreciation & amortization 2.07% 2.05% 2.23%
Operating income 5.63% 9.60% 8.99%
Interest - net 0.88% 0.92% 0.99%
Loss on extinguishment of debt 0.00% 0.68% 0.00%
Pre-tax earnings 4.76% 8.00% 8.00%
Income tax provision 1.51% 2.98% 3.24%
Net earnings 3.25% 5.02% 4.76%